OpenStudy (anonymous):

Find each percent increase or decrease to the nearest percent. from 30 to 24 A. Increased B. Decreased What is the nearest percentage

OpenStudy (shubhamsrg):

???

OpenStudy (hba):

i couldnt get ur question

OpenStudy (anonymous):

ITS TOO CONFUSING AHHHHHHHHH

ganeshie8 (ganeshie8):

which part you have problem with ?

OpenStudy (anonymous):

i'm having trouble with finding the percentage :/

ganeshie8 (ganeshie8):

ok. here is an easy process to find percentages for these : 30 --------> 24 lets call, initial value = 30 final value = 24

ganeshie8 (ganeshie8):

fine ? we are calling them so, because, we started with "30" and ending with "24" if it helps, think of them as, you got 30 marks in exam last time, now you got only 24

OpenStudy (anonymous):

okay

ganeshie8 (ganeshie8):

to find the percentage, do this : \(\huge \frac{\textbf{(final value) - (initial value)}}{\textbf{(initial value)}}*100\)

ganeshie8 (ganeshie8):

could you plugin above formula, initial value = 30 final value = 24

OpenStudy (anonymous):

:D

ganeshie8 (ganeshie8):

it becomes, \(\huge \frac{24 - 30}{20} * 100 = -30 \%\)

ganeshie8 (ganeshie8):

negative sign indicates it got Decreased !

ganeshie8 (ganeshie8):

any percentage you can calculate using above formula, just figure out initial value final value plugin hope it help... :)

OpenStudy (anonymous):

Yep and thank you :D

ganeshie8 (ganeshie8):

np.. yw :))